Undetermined Coefficients

4. y^ prime prime +2y^ prime +5y=1+e^ x



1
Expert's answer
2022-06-13T23:06:50-0400
"y''+2y'+5y=1+e^x"

The characteristic (auxiliary) equation for homogeneous differential equation is


"r^2+2r+5=0"

"(r+1)^2=-4"

"r=-1\pm 2i"

The general solution of the homogeneous differential equation is


"y_h=C_1e^{-x}\cos 2x+C_2e^{-x}\sin 2x"

Find a particular solution of the nonhomogeneous differential equation. 


"y_1=Ae^x+B"

"y_1'=Ae^x"

"y_1''=Ae^x"

Substitute


"Ae^x+2Ae^x+5Ae^x+5B=1+e^x"

"A=\dfrac{1}{8}, B=\dfrac{1}{5}"

"y_1=\dfrac{1}{8}e^x+\dfrac{1}{5}"

The general solution of the given differential equation is


"y_h=C_1e^{-x}\cos 2x+C_2e^{-x}\sin 2x+\dfrac{1}{8}e^x+\dfrac{1}{5}"